What us integral ydy /(x2+y2)power1/2 in trigometric form

Grade:11

1 Answers

Vikas TU
7 years ago
Let x = cosa = > dx/da= -sina …................(1)
     y = sina = > dy/da = cosa …...............(2)
From eqns. (1) and (2)
dx/dy = –sina/cosa..............................(3)

Now in integeral form, ydy /(x2+y2) = dxdx/dy =y/(x2 + y2)from eqns. (1), (2) and (3) we get,-sina/cosa = sinasina*(1 + 1/cosa) =0sina = 0   and cosa = -1a = +-npie  and a = (2n ± 1)pieTake the intersection.